# 22mm22mm25mm25mm

### v1

@@
Abstract.
This article studies canonical forms derived from the finest simultaneous block diagonalization of a set of symmetric matrices via congruence. Our technique relies on Harrison’s center theory, which is extended from a single higher degree form to multiple quadratic forms, hence a set of symmetric matrices. The algebraic structures of centers and the bijective relationship between the simultaneous block diagonalization via congruence and complete sets of orthogonal idempotents of centers are investigated. We provide an algorithm that mainly uses standard linear algebra tasks and several examples to demonstrate its effectiveness. In addition, this technique can be extended verbatim to the simultaneous block diagonalization of a set of Hermitian matrices via *-congruence.
Key words and phrases:
simultaneous block diagonalization via congruence, center
2020 Mathematics Subject Classification:
15A69, 15A20, 13P05
1. Introduction
1.1. Problem statement
@@
via congruence for all $1\leq i\leq m$, where $P\in GL_{n}(\mathbbm{k})$ is an invertible matrix, $B_{ij}\in\mathbbm{k}^{n_{j}\times n_{j}}$ and $\sum_{j=1}^{t}n_{j}=n$. This problem is called simultaneous block diagonalization via congruence (SBDC) of a set of symmetric matrices. If $t=n$, $A_{1},A_{2},\ldots,A_{m}$ are simultaneously diagonalized and it is referred to as simultaneous diagonalization via congruence (SDC). The SBDC of multiple symmetric matrices are classical problems in linear algebra and have been of interest to many researchers. By reformulating the original problem into one with less complexity, large-scale problems can be solved more efficiently. Thus, the SBDC has found applications in many areas, including quadratic programming, variational analysis and signal processing, see [3, 15, 23] and references therein.
1.2. Previous work
@@
As a special case, the SDC is less complex and has been studied thoroughly. Early work on the SDC of two real symmetric matrices was initialized in 1868 by Weierstrass [25] and pairs of quadratic forms were developed by authors like Finsler [7] and Calabi [4]. In 2007, Hiriart-Urruty [10] posed an open question about the SDC of a set of symmetric matrices. Until recently, Jiang and Li [15] extended the SDC to more than two real symmetric matrices, which was further investigated by [19]. Some papers, such as [1, 11, 16], considered the SDC of two Hermitian matrices and it was expanded to a set of Hermitian matrices in [17]. In more recent work [3], authors tackled the SDC for a set of general complex symmetric matrices. Additionally, some researchers explored the SDC using quadratic forms associated with symmetric matrices. The SDC of two quadratic forms has been approached in [2, 9, 26] and was extended to the almost SDC for a set of quadratic forms in [24].
@@
Compared to the SDC, which poses strict constraints on the matrices, the SBDC concerns the canonical forms of general sets of symmetric matrices and is much more difficult to solve. Uhlig [20] generalized the SDC of two real symmetric matrices to the SBDC of two real symmetric matrices based on their Jordan normal forms. His later papers [21, 22] further examined canonical pair forms of two real symmetric matrices and provided a new derivation. Until now, the SBDC of more than two symmetric matrices or Hermitian matrices has not been explored and remains a challenging problem.
1.3. Methods and results
Harrison [8] introduced centers for higher degree forms, which are highly effective for decomposing a higher degree form into a sum of forms in disjoint sets of variables [13, 14]. Recall that the center of a higher degree form, i.e. a homogeneous polynomial with degree $>2$, $f(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ is defined as
$Z(f):=\{X\in\mathbbm{k}^{n\times n}\mid(H_{f}X)^{T}=H_{f}X\},$
@@ 1.3. Methods and results
where $H_{f}$ is the Hessian matrix of $f$. It was shown that the direct sum decomposition of a form is in bijection with the decomposition of the unit of its center algebra into orthogonal idempotents. A criterion and an algorithm mostly based on standard linear algebra tasks were provided to decompose arbitrary forms. Harrison’s center theory was extended to the simultaneous direct sum decompositions of a set of not necessarily homogeneous polynomials in [6]. The center of a set of multivariate polynomials $f_{i}(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ for $1\leq i\leq m$ is defined as
[…]
It was proved that $Z(f_{1},f_{2},\dots,f_{m})$ is a special Jordan algebra and it was applied to the simultaneous direct sum decompositions of $f_{1},f_{2},\ldots,f_{m}$.
Due to the association between quadratic forms and symmetric matrices, we notice that the simultaneous direct sum decomposition of a set of quadratic forms is equivalent to the SBDC of a set of symmetric matrices. Thus, we extend Harrison’s center theory to the SBDC of a set of symmetric matrices. Centers of a set of quadratic forms defined in Equation (1.1) are reformulated to those of a set of symmetric matrices. The bijective relationship between the SBDC of a set of symmetric matrices and complete sets of orthogonal idempotents of their center is proven. A simple algorithm and several examples are provided. Finally, we generalize our approach for the SBDC of a set of symmetric matrices to the SBDC of a set of Hermitian matrices.
1.4. Organization of the paper
@@
The remainder of this paper is organized as follows. In Section 2 we introduce and apply centers to the SBDC of a set of symmetric matrices and give a simple algorithm. Section 3 presents more examples to demonstrate the effectiveness of this technique. Section 4 extends the center theory to the SBDC of a set of Hermitian matrices.
2. The center and simultaneous block diagoanlizations
2.1. The center of a set of symmetric matrices
@@
We begin by defining the center of a set of symmetric matrices. Let $A\in\mathbbm{k}^{n\times n}$ be a symmetric matrix and $f(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ be its associated quadratic form. Since $f$ can be represented as $f(x)=x^{T}Ax$, where $x=(x_{1},x_{2},\dots,x_{n})^{T}$, it is easy to work out $H_{f}=2A$. Thus, we can reformulate the center of a set of quadratic forms in Equation (1.1) to that of a set of symmetric matrices.
@@
Under the assumption that the characteristic of $\mathbbm{k}$ is not equal to $2$, a set of symmetric matrices and their associated quadratic forms share the same center. For clarity, we focus on symmetric matrices instead of quadratic forms in the rest of this article.
@@
Consider symmetric matrices $A_{1}=\begin{pmatrix}-2&2&-2\\ 2&2&0\\ -2&0&-1\end{pmatrix}$ and $A_{2}=\begin{pmatrix}5&7&-1\\ 7&5&1\\ -1&1&-1\end{pmatrix}$. According to Equation (2.1), we compute the center of $A_{1}$ and $A_{2}$ as
@@
Example 2.4.
Given the following symmetric matrices
$A_{1}=\begin{pmatrix}1&1&1\\ 1&1&0\\ 1&0&1\end{pmatrix},\quad A_{2}=\begin{pmatrix}-2&2&-2\\ 2&2&0\\ -2&0&-1\end{pmatrix},\quad A_{3}=\begin{pmatrix}1&3&-2\\ 3&3&0\\ -2&0&2\end{pmatrix}.$
Their center is calculated as
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}38a+22b&0&24a\\ -19a-11b+11c&22c&-12a\\ 22a&0&22b\end{pmatrix}\vline\,a,b,c\in\mathbbm{k}\right\}.$
2.2. Jordan algebras
The center of a set of symmetric matrices in Equation (2.1) may not have an associative algebraic structure as the center of a single higher degree form. However, we observe that it has a special Jordan algebraic structure as discussed below. Some necessary definitions of Jordan algebras are included and more are given in [18].
Definition 2.5.
A Jordan algebra is a non-associative algebra, which consists of a vector space $J$ with a bilinear product $X\odot Y$ that satisfies the following identities:
$X\odot Y=Y\odot X,\quad(X^{2}\odot Y)\odot X=X^{2}\odot(Y\odot X).$
Let $A$ be an associative algebra over $\mathbbm{k}$. We attain a Jordan algebra $A^{+}$ with a symmetrized product
$X\odot Y=\frac{1}{2}(XY+YX),$
which is commutative and satisfies the above identities. A special Jordan algebra is a Jordan algebra that is isomorphic to a subalgebra of some $A^{+}$.
An idempotent is an element $e\in J$ that satisfies $e\odot e=e$. Two idempotents $e_{1}$ and $e_{2}$ are orthogonal if $e_{1}\odot e_{2}=0$. In a unital Jordan algebra $J$, a set of idempotents $e_{1},e_{2},\dots,e_{s}$ are referred to as a complete set of orthogonal idempotents if
$e_{i}\odot e_{i}=e_{i},\,\forall i,\quad e_{i}\odot e_{j}=0,\,\forall i\neq j,\quad\sum_{i=1}^{s}e_{i}=1.$
Remark 2.6.
If $e_{1},e_{2},\dots,e_{s}$ are a complete set of orthogonal idempotents in a unital Jordan subalgebra $J$ of $({\mathbbm{k}^{n\times n}})^{+}$, they must satisfy the following properties:
$e_{i}^{2}=e_{i},\,\forall i,\quad e_{i}e_{j}=0,\,\forall i\neq j,\quad\sum_{i=1}^{s}e_{i}=I_{n}.$
We will briefly explain $e_{i}e_{j}=0$ and the other two properties are obvious. Since an idempotent $e_{i}$ must satisfy $e_{i}^{2}=e_{i},$ there exists an invertible matrix $P$ such that $P^{-1}e_{i}P=\begin{pmatrix}I&0\\ 0&0\end{pmatrix}.$ Let $P^{-1}e_{j}P$ be represented as a block matrix $\begin{pmatrix}X&Y\\ Z&W\end{pmatrix}$. Since $e_{i}\odot e_{j}=\frac{1}{2}(e_{i}e_{j}+e_{j}e_{i})=0$ for any two orthogonal idempotents $e_{i},e_{j}$, we can see that $P^{-1}e_{j}P=\begin{pmatrix}0&0\\ 0&W\end{pmatrix}$ and thus $e_{i}e_{j}=0$.
Theorem 2.7.
Let $A_{1},A_{2},\ldots,A_{m}\in\mathbbm{k}^{n\times n}$ be a set of symmetric matrices and $Z(A_{1}$, $A_{2}$, $\ldots$, $A_{m})$ be their center. Then $(Z(A_{1},A_{2},\ldots,A_{m}),\odot)$ is a special Jordan algebra.
Proof.
By definition, any $X,Y\in Z(A_{1},A_{2},\ldots,A_{m})$ need to satisfy $X\odot Y=\frac{1}{2}(XY+YX)$. We have
$A_{i}(X\odot Y)=A_{i}\frac{(XY+YX)}{2}=\frac{(Y^{T}X^{T}+X^{T}Y^{T})}{2}A_{i}=(X\odot Y)^{T}A_{i}=[A_{i}(X\odot Y)]^{T},$
for $1\leq i\leq m$. This implies that $(Z(A_{1},A_{2},\ldots,A_{m}),\odot)$ is closed under symmetric product $\odot$ and hence a Jordan subalgebra of ${(\mathbbm{k}^{n\times n})}^{+}$. ∎
2.3. Simultaneous block diagoanlization of symmetric matrices
Based on the algebraic structures of centers $Z(A_{1},A_{2},\ldots,A_{m})$ discussed in Subsection 2.2, we simultaneously block diagonalize a set of symmetric matrices via congruence.
@@ 2.3. Simultaneous block diagoanlization of symmetric matrices
    There is a one-to-one correspondence between simultaneous block diagonalizations of $A_{1}$, $A_{2}$, $\ldots$, $A_{m}$ via congruence and complete sets of orthogonal idempotents of $Z(A_{1},A_{2},\ldots,A_{m})$.
@@
    $A_{1},A_{2},\ldots,A_{m}$ are simultaneously diagonalizable via congruence if and only if $Z(A_{1}$, $A_{2},\ldots,A_{m})$ has a complete set of $n$ orthogonal idempotents.
@@
A proof of item (1) is provided, which implies items (2) and (3). We begin by examining how centers of a set of symmetric matrices vary under the congruence transformation. Let $B_{1},B_{2},\ldots,B_{m}\in\mathbbm{k}^{n\times n}$ be a set of matrices. Suppose there exists an invertible matrix $P\in\operatorname{GL}_{n}(\mathbbm{k})$ such that $B_{i}=P^{T}A_{i}P$ for $1\leq i\leq m$. According to Equation (2.1), we have $(B_{i}Y)^{T}=B_{i}Y$ for any $Y\in Z(B_{1},B_{2},\ldots,B_{m})$. It is ready to see that $(B_{i}Y)^{T}=Y^{T}P^{T}A_{i}P=P^{T}A_{i}PY$, which shows that $PYP^{-1}\in Z(A_{1},A_{2},\ldots,A_{m})$. Consequently, we have $Z(B_{1},B_{2},\ldots,B_{m})=P^{-1}Z(A_{1},A_{2},\ldots,A_{m})P.$
@@
for all $1\leq i\leq m$. Since $B_{i}$ is a block diagonal matrix and can be written as $B_{i}=B_{i1}\oplus B_{i2}\oplus\cdots\oplus B_{it}$, we see that their centers can be represented as $Z(B_{1},B_{2},\ldots,B_{m})=Z(B_{11},B_{21},\ldots,B_{m1})\oplus Z(B_{12},B_{22},\ldots,B_{m2})\oplus\cdots\oplus Z(B_{1t},B_{2t},\ldots,B_{mt})$. Thus, there exists a correspondence between each identity matrix $I_{j}$ of $Z(B_{1j},B_{2j},\ldots,B_{mj})$ and an idempotent element $e_{j}$ of $Z(B_{1},B_{2},\ldots,B_{m})$. Note that $e_{j}$ is a block diagonal matrix that is written as $e_{j}=0\oplus\cdots\oplus I_{j}\oplus\cdots\oplus 0$. Since the center of $A_{1},A_{2},\ldots,A_{m}$ can be represented as $Z(A_{1},A_{2},\dots,A_{m})=PZ(B_{1},B_{2},\dots,B_{m})P^{-1}$, it is obvious that $Pe_{1}P^{-1},Pe_{2}P^{-1}$, $\ldots$, $Pe_{t}P^{-1}$ are a complete set of orthogonal idempotents of $Z(A_{1},A_{2},\ldots,A_{m})$.
Conversely, suppose there exists a complete set of orthogonal idempotents $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ in $Z(A_{1},A_{2},\ldots,A_{m})$. We have $\epsilon_{i}^{2}=\epsilon_{i}$ for all $i$ and $\epsilon_{i}\epsilon_{j}=0$ for all $i\neq j$ according to Remark 2.6. Since the $\epsilon_{i}$’s are diagonalizable matrices and they pairwise commute, $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ are simultaneously diagoanalizable. Therefore, there exists an invertible matrix $P\in\operatorname{GL}_{n}(\mathbbm{k})$ such that $P^{-1}\epsilon_{j}P=e_{j}$, which satisfy $\sum_{j=1}^{t}e_{j}=I_{n}$. Let $B_{i}=P^{T}A_{i}P$ for all $1\leq i\leq m$ as defined above. We now have $B_{i}e_{j}=(B_{i}e_{j})^{T}=e_{j}^{T}B_{i}^{T}=e_{j}B_{i}$ for $1\leq i\leq m$ and $1\leq j\leq t$. We further deduce that the $B_{i}$’s are block diagonal matrices and can be written as
@@
for all $1\leq i\leq m$. Therefore, a simultaneous block diagonalization of $A_{1},A_{2},\ldots,A_{m}$ via congruence corresponds to the complete set of orthogonal idempotents $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$. ∎
2.4. Algorithm for simultaneous block diagonalization
We describe the procedures of the SBDC of a set of symmetric matrices in Algorithm 1. It is a recursive process that begins with computing the center of these symmetric matrices according to Definition 2.1. Using the base matrices of the center, a pair of orthogonal idempotents $\epsilon_{1}$ and $\epsilon_{2}$ is calculated. If no orthogonal idempotent is found, the symmetric matrices are not simultaneously block diagonalizable via congruence. Otherwise, an invertible matrix $P$ is computed according to Theorem 2.8 and these symmetric matrices are simultaneously block diagonalized via congruence as a set of block diagonal matrices, each with two diagonal block submatrices.

### v2

# Simultaneous block diagonalization of symmetric matrices via congruence

@@
Abstract
This article studies canonical forms derived from the finest simultaneous block diagonalization of a set of symmetric matrices via congruence. Our technique relies on the center theory of a set of multivariate polynomials, which we adapt for a set of symmetric matrices. We establish a bijective relationship between the simultaneous block diagonalizations of these matrices via congruence and complete sets of orthogonal idempotents of their centers. Based on this framework, we provide an algorithm relying primarily on standard linear algebra tasks. We also extend this technique to the simultaneous block diagonalizations of a set of Hermitian matrices via *-congruence. In addition, we investigate the simultaneous orthogonal block diagonalizations for real matrices and the simultaneous unitary block diagonalizations for Hermitian matrices. Several examples are provided to demonstrate its effectiveness.
Corresponding author: hualin.huang@hqu.edu.cn
Keywords: simultaneous diagonalization via congruence, center, symmetric matrices, Hermitian matrices
MSC: 15A69, 15A20, 13P05
1 Introduction
1.1 Problem statement
@@
via congruence for all $1\leq i\leq m$, where $P\in\operatorname{GL}_{n}(\mathbbm{k})$ is an invertible matrix, $B_{ij}\in\mathbbm{k}^{n_{j}\times n_{j}}$ and $\sum_{j=1}^{t}n_{j}=n$. This problem is called simultaneous block diagonalization via congruence (SBDC) of a set of symmetric matrices. If $t=n$, $A_{1},A_{2},\ldots,A_{m}$ are simultaneously diagonalized and it is referred to as simultaneous diagonalization via congruence (SDC). The SBDC of multiple symmetric matrices is a classical problem in linear algebra and has been of interest to many researchers. By reformulating the original problem into one with less complexity, large-scale problems may be solved more efficiently. Thus, the SBDC has found applications in many areas, including quadratic programming, variational analysis and signal processing, see [bustamante2020solving, jiang2016simultaneous, vollgraf2006quadratic] and references therein.
Let $q_{1},q_{2},\ldots,q_{m}\in\mathbbm{k}[x_{1},x_{2},\ldots,x_{n}]$ be a set of quadratic forms defined by $q_{i}(x)=x^{T}A_{i}x$. It is of interest to find a change of variables $x=Py$ with $P\in\operatorname{GL}_{n}(\mathbbm{k})$ such that
$q_{i}(Py)=q_{i1}(y_{1},\dots,y_{a_{1}})+q_{i2}(y_{a_{1}+1},\dots,y_{a_{2}})+\dots+q_{it}(y_{a_{t-1}+1},\dots,y_{n})$
for all $1\leq i\leq m$. This is called the simultaneous direct sum decomposition or simultaneous diagonalization if $t=n$ for $q_{1},q_{2},\ldots,q_{m}.$ Due to the well-established equivalence between symmetric matrices and quadratic forms, the SDC and SBDC of a set of symmetric matrices correspond directly to the simultaneous diagonalization and direct sum decompositions of the associated set of quadratic forms. Specifically, the problem is equivalent to determining a change of variables $x=Py$ such that
$q_{i}(Py)=(Py)^{T}A_{i}(Py)=y^{T}(P^{T}A_{i}P)y,$
where $P^{T}A_{i}P$ are block diagonal matrices, as defined above.
1.2 Previous work
@@
As a special case, the SDC is less complex and has been studied thoroughly. Early work on the SDC of two real symmetric matrices was initiated by Weierstrass [weierstrass1868zur] in 1868, and pairs of quadratic forms were developed by authors like Finsler [finsler1936uber] and Calabi [calabi1964linear]. In 2007, Hiriart-Urruty [hiriart2007potpourri] posed an open question about the SDC of a set of real symmetric matrices. Until recently, Jiang and Li [jiang2016simultaneous] extended the SDC to more than two real symmetric matrices, which was further investigated by [nguyen2020simultaneous]. Some papers, such as [auyeung1970necessary, hong1986reduction, lancaster2005canonical], considered the SDC of two Hermitian matrices, and it was expanded to a set of Hermitian matrices in [le2022simultaneous]. In more recent work [bustamante2020solving], authors tackled the SDC for a set of general complex symmetric matrices. Additionally, some researchers explored the SDC using quadratic forms associated with symmetric matrices. The SDC of two quadratic forms has been approached in [becker1980necessary, hestenes1940theorem, wonenburger1966simultaneous], and was extended to the almost SDC for a set of quadratic forms in [wang2024notion].
@@
Compared to the SDC, which poses strict constraints on the matrices, the SBDC concerns the canonical forms of general sets of symmetric matrices and is much more difficult to solve. Uhlig [uhlig1973simultaneous] generalized the SDC of two real symmetric matrices to the SBDC of two real symmetric matrices based on their Jordan normal forms. His later papers [uhlig1976canonical, uhlig1979recurring] further examined canonical pair forms of two real symmetric matrices and provided a new derivation. Until now, the SBDC of more than two symmetric matrices or Hermitian matrices has not been explored and remains a challenging problem. Furthermore, while canonical forms for a single matrix under restricted transformations like orthogonal congruence and unitary *-congruence are known [hong1989canonical, horn2009canonical], extending the SBDC to a set of matrices under these transformations presents an even more complex challenge.
In [fang2025simultaneous], the authors extended Harrison’s center theory [harrison1975grothendieck] and applied it to the simultaneous direct sum decompositions of a general set of multivariate polynomials. Such investigation for a single homogeneous polynomial has been carried out in [huang2021diagonalizable, huang2021centres]. The problem of simultaneously decomposing a set of polynomials is connected to the simultaneous Waring decompositions of multiple polynomials [carlini2003waring, tiels2013coupled]. Authors [decuyper2019decoupling, dreesen2018decoupling, dreesen2015decoupling] then developed tensor decomposition-based techniques to simultaneously decompose several multivariate polynomials into sums of univariate polynomials. On the one hand, if we focus on a set of quadratic forms, then the idea and method of [fang2025simultaneous] can be refined to the present SBDC problem. On the other hand, there is a surprising equivalence between the simultaneous direct sum decomposition of a general set of multivariate polynomials and the SBDC of a set of symmetric matrices. This is explained in the next subsection with necessary notations.
1.3 Methods and results
@@
It has been established in [fang2025simultaneous] that a simultaneous direct sum decomposition of several forms corresponds to an additive decomposition of the unit of their center algebra into orthogonal idempotents. A criterion and an algorithm, mostly based on standard linear algebra tasks, were provided to simultaneously decompose arbitrary forms. The center of a set of multivariate polynomials $f_{i}(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ for $1\leq i\leq m$ is defined as
$Z(f_{1},f_{2},\dots,f_{m}):=\{X\in\mathbbm{k}^{n\times n}\mid(H_{f_{i}}X)^{T}=H_{f_{i}}X,\ 1\leq i\leq m\},$ (1.1)
where $H_{f_{i}}$ is the Hessian matrix of $f_{i}$. It was proved that $Z(f_{1},f_{2},\dots,f_{m})$ is a special Jordan algebra, and it was applied to the simultaneous direct sum decompositions of $f_{1},f_{2},\ldots,f_{m}$. The main idea of this paper is to restrict the polynomials to a set of quadratic forms and apply the machinery of [fang2025simultaneous] to the SBDC problem to derive more precise results.
As a critical foundation, the center of a set of quadratic forms defined in Equation (1.1) is reformulated into that of a set of symmetric matrices. We prove a bijective relationship between the SBDC of a set of symmetric matrices and complete sets of orthogonal idempotent elements of their center. Furthermore, we also address simultaneous orthogonal block diagonalization for real symmetric matrices. A simple algorithm and several examples are provided. In addition, we generalize our approach to the SBDC of a set of Hermitian matrices, including the special case of simultaneous unitary block diagonalization. It is worthy of remarking that the center of a set of multivariate polynomials is essentially the center of a set of symmetric matrices, and thus direct sum decompositions of arbitrary polynomials amount to the SBDC problem of symmetric matrices. Indeed, if we write the Hessian matrix of $f_{i}$ as a sum of scalar symmetric matrices multiplied by suitable monomials, then the equation in (1.1) can be rewritten as the one in (2.1), see Definition 2.1. Therefore, the SBDC problem includes many important problems as special cases, and our approach provides a highly unified solution to all of them.
1.4 Organization of the paper
@@
The remainder of this paper is organized as follows. In Section 2 we introduce and apply centers to the SBDC and orthogonal SBDC of a set of symmetric matrices, and give a simple algorithm. Section 3 presents more examples to demonstrate the effectiveness of this technique. Section 4 extends the center theory to the SBDC and the unitary SBDC of a set of Hermitian matrices.
2 The center and simultaneous block diagonalizations
2.1 The center of a set of symmetric matrices
@@
We begin by defining the center of a set of symmetric matrices. Let $A\in\mathbbm{k}^{n\times n}$ be a symmetric matrix and $q(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ be its associated quadratic form. Since $q$ can be represented as $q(x)=x^{T}Ax$, where $x=(x_{1},x_{2},\dots,x_{n})^{T}$, one readily finds that $H_{q}=2A$. Thus, we can reformulate the center of a set of quadratic forms in Equation (1.1) to the center of a set of symmetric matrices.
@@
Under the assumption that the characteristic of $\mathbbm{k}$ is not equal to $2$, a set of symmetric matrices and their associated quadratic forms share the same center. For clarity, we focus on symmetric matrices instead of quadratic forms in the rest of this article. It is obvious that $Z(A_{1},A_{2},\dots,A_{m})$ contains all scalar matrices.
@@
Consider the pair of symmetric matrices $A_{1}=\begin{pmatrix}-2&2&-2\\ 2&2&0\\ -2&0&-1\end{pmatrix}$ and $A_{2}=\begin{pmatrix}5&7&-1\\ 7&5&1\\ -1&1&-1\end{pmatrix}$. According to Equation (2.1), we compute the center of $A_{1}$ and $A_{2}$ as
@@
The center of a set of symmetric matrices may not have an associative algebraic structure as the center of a single higher degree form. However, it has a special Jordan algebraic structure as observed in [fang2025simultaneous]. As a matter of fact, if $X,Y\in Z(A_{1},A_{2},\dots,A_{m})$, then $X\odot Y:=\frac{1}{2}(XY+YX)\in Z(A_{1},A_{2},\dots,A_{m})$, hence $(Z(A_{1},A_{2},\dots,A_{m}),\odot)$ becomes a Jordan subalgebra of the full matrix Jordan algebra under the symmetric matrix product $\odot$. It is not hard to see that, a complete set of orthogonal idempotents of $(Z(A_{1},A_{2},\dots,A_{m}),\odot)$ is a set of mutually orthogonal matrices in $Z(A_{1},A_{2},\dots,A_{m})$ which sum up to the identity matrix, see [fang2025simultaneous] for more details.
2.2 Simultaneous block diagonalization of symmetric matrices
We call $\epsilon_{1},\epsilon_{2},\dots,\epsilon_{t}\in Z(A_{1},A_{2},\ldots,A_{m})$ a complete set of orthogonal idempotent elements if
$\epsilon_{i}^{2}=\epsilon_{i}\ \forall i,\quad\epsilon_{i}\epsilon_{j}=0\ \forall i\neq j,\quad\sum_{i}\epsilon_{i}=I_{n}.$
It turns out that complete sets of orthogonal idempotent elements of centers are crux for the purpose of simultaneously block diagonalizing a set of symmetric matrices via congruence.
@@ 2.2 Simultaneous block diagonalization of symmetric matrices
    There is a one-to-one correspondence between simultaneous block diagonalizations of $A_{1},A_{2},\ldots,A_{m}$ via congruence and complete sets of orthogonal idempotent elements of $Z(A_{1},A_{2},\ldots,A_{m})$.
@@
    $A_{1},A_{2},\ldots,A_{m}$ are simultaneously diagonalizable via congruence if and only if $Z(A_{1},A_{2},\ldots,A_{m})$ has a complete set of $n$ orthogonal idempotent elements.
  • (4)
    The finest simultaneous block diagonalization for $A_{1},A_{2},\ldots,A_{m}$ via congruence is unique up to equivalence.
@@
A proof of item (1) is provided, which evidently implies items (2) and (3). Item (4) is a direct consequence of the uniqueness of the complete set of primitive orthogonal idempotent elements of unital Jordan algebras up to equivalence, see [mccrimmon2003taste] for more details.
We begin by examining how centers of a set of symmetric matrices vary under the congruence transformation. Let $B_{1},B_{2},\ldots,B_{m}\in\mathbbm{k}^{n\times n}$ be a set of symmetric matrices. Suppose there exists an invertible matrix $P\in\operatorname{GL}_{n}(\mathbbm{k})$ such that $B_{i}=P^{T}A_{i}P$ for $1\leq i\leq m$. According to Equation (2.1), we have $(B_{i}Y)^{T}=B_{i}Y$ for any $Y\in Z(B_{1},B_{2},\ldots,B_{m})$. It follows that $(B_{i}Y)^{T}=Y^{T}P^{T}A_{i}P=P^{T}A_{i}PY$, which shows that $PYP^{-1}\in Z(A_{1},A_{2},\ldots,A_{m})$. Consequently, we have $Z(B_{1},B_{2},\ldots,B_{m})=P^{-1}Z(A_{1},A_{2},\ldots,A_{m})P.$
@@
for all $1\leq i\leq m$. For all $1\leq j\leq m$, let
$e_{j}=\begin{pmatrix}0&&&&\\ &\ddots&&&\\ &&I_{j}&&\\ &&&\ddots&\\ &&&&0\end{pmatrix}$
be block diagonal matrices which are conformal with the $B_{i}$’s. It is clear that the $e_{j}$’s are a complete set of orthogonal idempotent elements of $Z(B_{1},B_{2},\dots,B_{m})$. Since $Z(A_{1},A_{2},\dots,A_{m})=PZ(B_{1},B_{2},\dots,B_{m})P^{-1}$, it is obvious that $Pe_{1}P^{-1},Pe_{2}P^{-1},\ldots,Pe_{t}P^{-1}$ are a complete set of orthogonal idempotent elements of $Z(A_{1},A_{2},\ldots,A_{m})$.
Conversely, suppose $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ in $Z(A_{1},A_{2},\ldots,A_{m})$ are a complete set of orthogonal idempotent elements. Since the $\epsilon_{i}$’s are diagonalizable matrices and they pairwise commute, $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ are simultaneously diagonalizable. Therefore, there exists an invertible matrix $P\in\operatorname{GL}_{n}(\mathbbm{k})$ such that $P^{-1}\epsilon_{j}P=e_{j}$ for all $1\leq j\leq t$ as above. Let $B_{i}=P^{T}A_{i}P$ for all $1\leq i\leq m$. We now have $B_{i}e_{j}=(B_{i}e_{j})^{T}=e_{j}^{T}B_{i}^{T}=e_{j}B_{i}$ for $1\leq i\leq m$ and $1\leq j\leq t$. We further deduce that the $B_{i}$’s are block diagonal matrices and can be written as
@@
for all $1\leq i\leq m$. Therefore, a simultaneous block diagonalization of $A_{1},A_{2},\ldots,A_{m}$ via congruence corresponds to the complete set of orthogonal idempotent elements $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$. ∎
For real symmetric matrices, it is more appropriate to consider simultaneous congruence via orthogonal matrices in some specific problems. Our approach can be easily adapted to this situation.
Corollary 2.5 (Orthogonal congruence).
Suppose $A_{1},A_{2},\ldots,A_{m}\in\mathbbm{R}^{n\times n}$ are a set of real symmetric matrices. Then there is a one-to-one correspondence between simultaneous block diagonalizations of $A_{1},A_{2},\ldots,A_{m}$ via orthogonal congruence and complete sets of symmetric orthogonal idempotent elements of $Z(A_{1},A_{2},\ldots,A_{m})$.
Proof.
The correspondence for general cases was established in Theorem 2.4. We now refine it to the real orthogonal case. Suppose $Q\in\operatorname{O}_{n}(\mathbbm{R})$ is an invertible orthogonal matrix and $B_{i}=Q^{T}A_{i}Q$ are block diagonal as defined above. Then the corresponding idempotent elements $e_{1},e_{2},\dots,e_{t}$ of $Z(B_{1},B_{2},\ldots,B_{m})$ are real, symmetric, and mutually orthogonal. Since $Q^{-1}=Q^{T}$, we have $Z(A_{1},A_{2},\ldots,A_{m})=QZ(B_{1},B_{2},\ldots,B_{m})Q^{T}$. Thus, $Qe_{1}Q^{T},Qe_{2}Q^{T},\ldots,$ $Qe_{t}Q^{T}$ form a complete set of symmetric orthogonal idempotent elements of $Z(A_{1},A_{2},\ldots,A_{m})$.
Conversely, assume $Z(A_{1},A_{2}\dots,A_{m})$ contains a complete set of symmetric orthogonal idempotent elements $\epsilon_{1},\epsilon_{2},\dots,\epsilon_{t}$. Since $\epsilon_{1},\epsilon_{2},\dots,\epsilon_{t}$ are real, symmetric, and pairwise commute, they are simultaneously diagonalizable by an orthogonal matrix $Q\in\operatorname{O}_{n}(\mathbbm{R})$ such that $Q^{T}\epsilon_{j}Q=e_{j}$. The same argument as in Theorem 2.4 shows that $Q^{T}A_{i}Q$ commutes with all $e_{j}$ and is therefore block diagonal. Hence $Q$ simultaneously block diagonalizes $A_{1},A_{2},\dots,A_{m}$ via orthogonal congruence. This completes the correspondence. ∎
2.3 Algorithm for simultaneous block diagonalization
We describe the procedures for the SBDC of a set of symmetric matrices in Algorithm 1. It is a recursive process that begins with computing the center of these symmetric matrices according to Definition 2.1. Using the base matrices of the center, a nontrivial idempotent $\epsilon_{1}$ is found, and the second orthogonal idempotent is then defined as $\epsilon_{2}=I-\epsilon_{1}$. If no orthogonal idempotent is found, the symmetric matrices are not simultaneously block diagonalizable via congruence. Otherwise, an invertible matrix $P$ is computed according to Theorem 2.4 and these symmetric matrices are simultaneously block diagonalized via congruence as a set of block diagonal matrices, each with two diagonal block submatrices.